# Pointwise convergence of ergodic averages with Möbius weight

### v1

@@ 6. Proofs of the pointwise ergodic theorems
(6.8) $\begin{split}&\left|\frac{1}{N^{d+1}}\sum_{m\in[N^{d}],n\in[N]}\frac{\theta(n)}{(\log N)^{K}}\phi(T^{m}x)f_{1}(T^{m+P_{1}(n)}x)\cdots f_{k}(T^{m+P_{k}(n)}x)\right|\\ &\quad\ll_{A,P_{1},\ldots,P_{k}}\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log N)^{-2-K-A}.\end{split}$
For proving this estimate, we may assume that $\phi,f_{1},\ldots,f_{k}$ are supported on $[-CN^{,}CN^{d}]$ for some $C\ll_{a,P}1$. Note also that by assumption for $n\in[N]$ we have $|\theta(n)|/(\log N)^{K}\leq 1$.
If $d=1$ (that is, all the $P_{i}$ of degree at most $1$), the estimate (6.8) follows by combining Lemma 4.6 and the assumption (6.2). If instead $d\geq 2$ (in which case by assumption the $P_{i}$ have distinct degrees), (6.8) follows by combining Theorem 4.1 and the assumption (6.2). ∎
We then turn to Theorem 1.4. This theorem is an immediate consequence of the following general result combined with Lemma 5.1(ii).
Theorem 6.4 (Pointwise convergence of double-logarithmic polynomial ergodic averages with nice weight).
Let $K\in\mathbb{N}$. Let $\theta\colon\mathbb{N}\to\mathbb{C}$ satisfy $|\theta(n)|\leq K\log n$ for all $n\in\mathbb{N}$, and suppose that there exist some infinite sets $S_{1},\ldots,S_{K}\subset\mathbb{N}$ such that, for all large enough $N\geq 1$ and any $n\in[N]$, we have
(6.9) $|\theta(n)|\leq\sum_{j=1}^{K}\frac{N}{|S_{j}\cap[N]|}1_{S_{j}}(n).$
Suppose also that for all $s\in\mathbb{N}$ and $M\geq 3$ we have
(6.10) $\|\theta\|_{U^{s}[M]}\ll_{s}(\log\log M)^{-\gamma_{s}}$
for some $\gamma_{s}>0$.
Let $(X,\nu,T)$ be a measure-preserving system. Let $k\in\mathbb{N}$, let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients, and let $1<q_{1},\ldots,q_{k}\leq\infty$ satisfy $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}<1$. Then, there exists $c=c(P_{1},\ldots,P_{k},q_{1},\ldots,q_{k})>0$ such that, for any $f_{1}\in L^{q_{1}}(X),\ldots,f_{k}\in L^{q_{k}}(X)$ and $N\geq 3$, we have
(6.11) $\lim_{N\to\infty}\,(\log\log N)^{c}\mathbb{E}_{n\leq N}^{\log\log}\theta(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)=0$
for almost all $x\in X$.
Theorem 6.1 will also be reduced to an $L^{1}$ estimate.
Proposition 6.5.
$\|\mathbb{E}_{n\leq N}^{\log\log}\theta(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)\|_{L^{1}(X)}$
$\quad\ll_{K,P_{1},\ldots,P_{k}}((1+\|f_{1}\|_{L^{q_{1}}(X)})\cdots(1+\|f_{k}\|_{L^{q_{k}}(X)}))^{B(q_{1},\ldots,q_{k})}(\log\log N)^{-c^{\prime}}$
for some $c=c(P_{1},\ldots,P_{k})>0$ and $B(q_{1},\ldots,q_{k})>0$.
Proof of Theorem 6.4 assuming Proposition 6.5.
We are going to apply Lemma 5.4(ii). Note that for any bounded sequence $(a(n))_{n\in\mathbb{N}}$ and any $M\geq N$ we have
$\left|\frac{1}{\log\log M}\sum_{n\leq M}\frac{a(n)}{n\log n}-\frac{1}{\log\log N}\sum_{n\leq N}\frac{a(n)}{n\log n}\right|$
$\quad\ll\left(\frac{1}{\log\log N}-\frac{1}{\log\log M}\right)\sum_{n\leq M}\frac{|a(n)|}{n\log n}+\frac{1}{\log\log M}\sum_{N\leq n\leq M}\frac{|a(n)|}{n\log n}$
$\quad\ll\frac{\log\log M-\log\log N+1}{\log\log N}\frac{1}{\log\log M}\sum_{\ell\leq(\log M)/(\log 2)+1}\frac{1}{\ell}\mathbb{E}_{n\leq 2^{\ell}}|a(n)|$
$\quad\quad+\frac{1}{\log\log M}\sum_{(\log N)/(\log 2)\leq\ell\leq(\log M)/(\log 2)+1}\frac{1}{\ell}\mathbb{E}_{n\leq 2^{\ell}}|a(n)|.$
Using this with $a(n)=\theta(n)\prod_{j\in[k]}f_{j}(T^{P_{j}(n)}x)$, $M\in[N,\exp((\log N)^{1+(\log\log N)^{-c}})]$, and using Lemma 5.5 and the assumption (6.9), we obtain
$\|\mathbb{E}_{n\leq M}^{\log\log}\theta(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)-\mathbb{E}_{n\leq N}^{\log\log}\theta(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)\|_{L^{1}(X)}$
$\quad\ll(\log\log N)^{-c}\max_{j\in[K]}\max_{\ell\leq(\log M)/(\log 2)+1}\left(\mathbb{E}_{n\in S_{j}\cap[2^{\ell}]}\int_{X}|f_{1}(T^{P_{1}(n)}x)|\cdots|f_{k}(T^{P_{k}(n)}x)|\textnormal{ d}\nu(x)\right)$
$\quad\ll(\log\log N)^{-c}\|f_{1}\|_{L^{q_{1}}(X)}\cdots\|f_{k}\|_{L^{q_{k}}(X)}.$
Hence, from Lemma 5.4(ii) it follows that Proposition 6.5 implies Theorem 1.2. ∎
Proof of Proposition 6.5.
We again first reduce to the bounded case.
Reduction to the case of bounded functions. Suppose that the case $q_{1}=\ldots=q_{k}=\infty$ has been proven. Let $1<q_{1},\ldots,q_{k}<\infty$, $f_{i}\in L^{q_{i}}(X)$, and let $\kappa>0$ be small enough in terms of $P_{1},\ldots,P_{k}$. For $i\in\{1,\ldots,k\}$, we split
$f_{i}=f_{i,N}^{(1)}+f_{i,N}^{(2)},\quad\textnormal{where}\quad f_{i,N}^{(1)}(x)=f_{i}(x)1_{|f_{i}(x)|\leq(\log\log N)^{\kappa}},\,\,f_{i,N}^{(2)}(x)=f_{i}(x)1_{|f_{i}(x)|>(\log\log N)^{\kappa}}.$
Then, denoting by $\widetilde{A}_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1},\ldots,f_{k})$ the function inside the norm in (6.11), by linearity we obtain
$\widetilde{A}_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1},\ldots,f_{k})$ $=\widetilde{A}_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1,N}^{(1)},\ldots,f_{k,N}^{(1)})$
$\quad+\sum_{(i_{1},\ldots,i_{k})\in[2]^{k}\setminus(1,\ldots,1)}\widetilde{A}_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1,N}^{(i_{1})},\ldots,f_{k,N}^{(i_{k})}).$
Recalling that $\kappa>0$ is small enough, by the case of bounded functions we have
$\|\widetilde{A}_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1,N}^{(1)},\ldots,f_{k,N}^{(1)})\|_{L^{1}(X)}\ll_{k,P_{1},\ldots,P_{k},\kappa}(\log\log N)^{-c/2},$
where $c=c(P_{1},\ldots,P_{k})$ is as in (6.11) in the case of bounded functions.
Note that for any $q^{\prime}<q$ and $i\in\{1,\ldots,k\}$ we have
(6.12) $\int_{X}|f_{i,N}^{(2)}(x)|^{q^{\prime}}\textnormal{ d}\nu(x)\leq(\log\log N)^{-\kappa\cdot(q^{\prime}-q)}\int_{X}|f_{i}(x)|^{q}\textnormal{ d}\nu(x).$
Fix some $q_{i}^{\prime}\in(1,q_{i})$ with $\frac{1}{q_{1}^{\prime}}+\ldots+\frac{1}{q_{k}^{\prime}}=1$. Note also that, by the assumption on $\theta$, for $n\in[N]$ we have
$|\theta(n)|\leq\theta_{1}(n)+\cdots+\theta_{K}(n),\quad\textnormal{where}\quad\theta_{j}(n)=\frac{N}{|S_{j}\cap[N]|}1_{S_{j}}(n).$
Now, by Lemma 5.5 and (6.12), for any $(i_{1},\ldots,i_{k})\neq(1,\ldots,1)$ we get
$\|\widetilde{A}_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1,N}^{(i_{1})},\ldots,f_{k,N}^{(i_{k})})\|_{L^{1}(X)}$
$\quad\ll_{K}\max_{1\leq j\leq J}\|\widetilde{A}_{N}^{P_{1},\ldots,P_{k}}(\theta_{j};f_{1,N}^{(i_{1})},\ldots,f_{k,N}^{(i_{k})})\|_{L^{1}(X)}$
$\quad\leq(\log\log N)^{-c^{\prime}\kappa}\|f_{1}\|_{L^{q_{1}}(X)}\cdots\|f_{k}\|_{L^{q_{k}}(X)},$
where $c^{\prime}=\frac{q_{1}}{q_{1}^{\prime}}+\cdots+\frac{q_{k}}{q_{k}^{\prime}}-k>0$.
The case of bounded functions. Let $P_{1},\ldots,P_{k}$ be fixed polynomials with integer coefficients and with highest degree $d\in\mathbb{N}$. Let $(X,\nu,T)$ be a measure-preserving system, and let $f_{1},\ldots,f_{k}\in L^{\infty}(X)$. In the rest of the proof, we abbreviate $\widetilde{A}_{N}(x)\coloneqq\widetilde{A}_{N}(\theta;f_{1},\ldots,f_{k})(x)$.
By Cauchy–Schwarz, it suffices to show that
$\int_{X}|\widetilde{A}_{N}(x)|^{2}\textnormal{ d}\nu(x)\ll\|f_{1}\|_{L^{\infty}(X)}^{2}\cdots\|f_{k}\|_{L^{\infty}(X)}^{2}(\log\log N)^{-c}.$
From (6.9) we have $\limsup_{N\to\infty}\|\widetilde{A}_{N}\|_{L^{\infty}(X)}\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}$, so it suffices to show that for any $\phi\in L^{\infty}(X)$ we have
$\int_{X}\phi(x)\widetilde{A}_{N}(x)\textnormal{ d}\nu(x)\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log\log N)^{-c}.$
This can be written out as
$\int_{X}\mathbb{E}_{n\leq N}^{\log\log}\theta(n)\phi(x)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)\textnormal{ d}\nu(x)$
$\quad\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log\log N)^{-c}.$
Let $N_{0}=\exp(\exp((\log\log N)^{1/2}))$. By partial summation and the assumptions (6.9) and $|\theta(n)|\leq K\log n$, we have the bound
$\sum_{n\leq N_{0}}\frac{|\theta(n)|}{n\log n}\ll_{K}(\log\log N)^{1/2},$
so it suffices to show that
$\int_{X}\frac{1}{\log\log N}\sum_{N_{0}\leq n\leq N}\frac{\theta(n)}{n\log n}\phi(x)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)\textnormal{d}\nu(x)$
$\quad\ll(\log\log N)^{-c}\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}.$
By partial summation and the assumptions on $\theta$, this follows if we show uniformly for $N^{\prime}\in[N_{0},N]$ the bound
$\int_{X}\frac{1}{N^{\prime}}\sum_{n\in[N^{\prime}]}\theta(n)\phi(x)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)\textnormal{ d}\nu(x)$
$\quad\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log\log N)^{-c}.$
By the $T$-invariance of $\nu$, it suffices to show that
(6.13) $\begin{split}&\int_{X}\frac{1}{(N^{\prime})^{d+1}}\sum_{m\in[(N^{\prime})^{d}],n\in[N^{\prime}]}\theta(n)\phi(T^{m}x)f_{1}(T^{m+P_{1}(n)}x)\cdots f_{k}(T^{m+P_{k}(n)}x)\textnormal{ d}\nu(x)\\ &\quad\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log\log N)^{-c}.\end{split}$
We can restrict the integral in (6.13) to the full measure set $X^{\prime}$ where $\|\phi(x)\|\leq\|\phi\|_{L^{\infty}(X)}$ and $|f_{i}(x)|\leq\|f_{i}\|_{L^{\infty}(X)}$ for all $i\in[k]$. Combining Theorem 4.2 and the assumption (6.10) (and assuming as we may that $\phi,f_{1},\ldots,f_{k}$ are supported on $[-CN^{d},CN^{d}]$ for some $C\ll 1$), we see that there is some $\gamma=\gamma(P_{1},\ldots,P_{k})>0$ such that for all $x\in X^{\prime}$ we have the bound
$\left|\frac{1}{(N^{\prime})^{d+1}}\sum_{m\in[(N^{\prime})^{d}],n\in[N^{\prime}]}\theta(n)\phi(T^{m}x)f_{1}(T^{m+P_{1}(n)}x)\cdots f_{k}(T^{m+P_{k}(n)}x)\right|$
$\quad\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log\log N^{\prime})^{-\gamma}$
$\quad\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log\log N)^{-\gamma/2}.$
Now the desired bound (6.13) follows with $c=\gamma/2$. ∎
Finally, we shall prove Theorem 1.3. This follows from the next theorem and Lemma 5.1(ii).
Theorem 6.6 (Pointwise convergence of polynomial ergodic averages with nice weight along a subsequence).
Let $K\in\mathbb{N}$. Let $\theta\colon\mathbb{N}\to\mathbb{C}$ and suppose that there exist some infinite sets $S_{1},\ldots,S_{K}\subset\mathbb{N}$ such that, for all large enough $N\geq 1$ and any $n\in[N]$, we have
(6.14) $|\theta(n)|\leq\sum_{j=1}^{K}\frac{N}{|S_{j}\cap[N]|}1_{S_{j}}(n).$
Suppose also that for all $s\in\mathbb{N}$ and $M\geq 3$ we have
(6.15) $\|\theta\|_{U^{s}[M]}\ll_{s}(\log\log M)^{-\gamma_{s}}$
for some $\gamma_{s}>0$.
Let $N_{j}\gg\exp(\exp(\exp((\log j)^{2}))$ for $j\in\mathbb{N}$. Let $(X,\nu,T)$ be a measure-preserving system. Let $k\in\mathbb{N}$, let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients, and let $1<q_{1},\ldots,q_{k}\leq\infty$ satisfy $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}<1$. Then, for any $f_{1}\in L^{q_{1}}(X),\ldots,f_{k}\in L^{q_{k}}(X)$ and $j\geq 1$, we have
$\lim_{j\to\infty}\frac{\exp((\log j)^{3/2})}{N_{j}}\sum_{n\leq N_{j}}\theta(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)=0$
for almost all $x\in X$.
This is a direct consequence of the following $L^{1}$ estimate.
Proposition 6.7.
(6.16) $\begin{split}&\|\mathbb{E}_{n\leq N}\theta(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)\|_{L^{1}(X)}\\ &\quad\ll_{A,K,P_{1},\ldots,P_{k}}((1+\|f_{1}\|_{L^{q_{1}}(X)})\cdots(1+\|f_{k}\|_{L^{q_{k}}(X)}))^{B(q_{1},\ldots,q_{k})}(\log\log N)^{-c}\end{split}$
for some $c=c(P_{1},\ldots,P_{k})>0$ and $B(q_{1},\ldots,q_{k})>0$.
Proof of Theorem 6.6 assuming Proposition 6.7.
This is immediate from Lemma 5.4(iii). ∎
Proof of Proposition 6.7.
By the same argument as in the proof of Proposition 6.5, we may reduce matters to the case of bounded functions.
Assume now that $q_{1}=\ldots=q_{k}=\infty$. Let $P_{1},\ldots,P_{k}$ be fixed polynomials with integer coefficients with highest degree $d\in\mathbb{N}$. Let $(X,\nu,T)$ be a measure-preserving system, and let $f_{1},\ldots,f_{k}\in L^{\infty}(X)$. We write $A_{N}(x)\coloneqq\mathbb{E}_{n\leq N}\theta(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)$.
By Cauchy–Schwarz, it suffices to show that
$\|A_{N}\|_{L^{2}(X)}\ll\|f_{1}\|_{L^{\infty}(X)}^{2}\cdots\|f_{k}\|_{L^{\infty}(X)}^{2}(\log\log N)^{-c}$
for some $c=c(P_{1},\ldots,P_{k})>0$.
The assumption (6.14) and Lemma 5.5 give
$\|A_{N}\|_{L^{\infty}(X)}\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)},$
so it suffices to show that
$\int_{X}\phi(x)A_{N}(x)\textnormal{ d}\nu(x)\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log\log N)^{-c}$
for all $\phi\in L^{\infty}(X)$ and $N\geq 3$. By the $T$-invariance of $\nu$, this is equivalent to
$\int_{X}\frac{1}{N^{d+1}}\sum_{m\in[N^{d}],n\in[N]}\theta(n)\phi(T^{m}x)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)\textnormal{ d}\nu(x)$
$\quad\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log\log N)^{-c}$
for all $\phi\in L^{\infty}(X)$. But this estimate follows from Theorem 4.2 combined with the assumption (6.15). ∎
@@
(6.17) $\begin{split}&\left|\frac{1}{N^{4}}\sum_{m_{1},m_{2},m_{3},n\in[N]}\mu(n)\phi(T_{1}^{m_{1}}T_{2}^{m_{2}}T_{3}^{m_{3}}x)f_{1}(T_{1}^{m_{1}+n}T_{2}^{m_{2}}T_{3}^{m_{3}}x)\cdots f_{3}(T_{1}^{m_{1}}T_{2}^{m_{2}}T_{3}^{m_{3}+n}x)\right|\\ &\quad\ll_{A}\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{3}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log N)^{-A}\end{split}$
for any functions $\phi,f_{1},f_{2},f_{3}\in L^{\infty}(X)$ and any $x\in X$ for which $|\phi(x)|\leq\|\phi\|_{L^{\infty}(X)},|f_{j}|\leq\|f_{j}\|_{L^{\infty}(X)}$. By setting
$F_{j}((m_{1},m_{2},m_{3}))$ $=f_{j}(T_{1}^{m_{1}}T_{2}^{m_{2}}T_{3}^{m_{3}}x),$
$G((m_{1},m_{2},m_{3}))$ $=\phi(T_{1}^{m_{1}}T_{2}^{m_{2}}T_{3}^{m_{3}}x),$
the estimate (6.17) can be rewritten as
$\left|\frac{1}{N^{4}}\sum_{\mathbf{m}\in[N]^{3},n\in[N]}\mu(n)G(\mathbf{m})F_{1}(\mathbf{m}+(n,0,0))F_{2}(\mathbf{m}+(0,n,0))F_{3}(\mathbf{m}+(0,0,n))\right|$
$\quad\ll_{A}\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{3}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log N)^{-A}.$
But this follows directly from Lemma 4.7. ∎

### v2

@@ 7. Proofs of the pointwise ergodic theorems
(7.8) $\begin{split}&\left|\frac{1}{N^{d+1}}\sum_{m\in[N^{d}],n\in[N]}\frac{\theta(n)}{(\log N)^{A}}\phi(T^{m}x)f_{1}(T^{m+P_{1}(n)}x)\cdots f_{k}(T^{m+P_{k}(n)}x)\right|\\ &\quad\ll_{A,P_{1},\ldots,P_{k}}\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log N)^{-2A}.\end{split}$
Write $\theta=\theta_{1}+\theta_{2}$ where $\theta_{1}(n)=\theta(n)1_{|\theta(n)|\leq(\log n)^{A}}$, $\theta_{2}(n)=\theta(n)1_{|\theta(n)|>(\log n)^{A}}$. Since $d(n)^{K}\geq(\log n)^{A-K}$ in the support of $\theta_{2}$, we have
[…]
if $A$ is large enough in terms of $K$.
Now, if assumption (i) of the theorem holds, by the triangle inequality and (7.9) we have
[…]
If instead assumption (ii) holds, we similarly have
(7.11) $\|\theta_{1}\|_{u^{s}[N]}\ll_{A,s,P_{1},\ldots,P_{k}}(\log N)^{-(A\log\log A)/2}.$
In either case, since we may assume that $\phi,f_{1},\ldots,f_{k}$ are supported on $[-CN^{,}CN^{d}]$ for some $C\ll_{P}1$ and since $|\theta_{1}(n)|(\log n)^{-A}\leq 1$, we can use either Theorem 4.2 or 4.1 (depending on whether we have (7.11) or (7.10)) to conclude that (7.8) holds with $\theta_{1}$ in place of $\theta$. Hence it suffices to show that (7.8) holds with $\theta_{2}$ in place of $\theta$. For this it suffices to show that
$\mathbb{E}_{n\leq N}|\theta_{2}(n)|\ll(\log N)^{-A}.$
Here the left-hand side is
$\leq(\log N)^{K-2(A-K)}\mathbb{E}_{n\leq N}d(n)^{2K}\ll(\log N)^{-A}$
if $A$ is large enough in terms of $K$. This completes the proof. ∎
@@
(7.12) $\begin{split}&\left|\frac{1}{N^{k+1}}\sum_{m_{1},\ldots,m_{k},n\in[N]}\mu(n)\phi(T_{1}^{m_{1}}\cdots T_{k}^{m_{k}}x)\prod_{j=1}^{k}f_{j}(T_{1}^{m_{1}}\cdots T_{k}^{m_{k}}T_{j}^{n}x)\right|\\ &\quad\ll_{A}\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log N)^{-A}\end{split}$
for any functions $\phi,f_{1},\ldots,f_{k}\in L^{\infty}(X)$ and any $x\in X$ for which $|\phi(x)|\leq\|\phi\|_{L^{\infty}(X)},|f_{j}|\leq\|f_{j}\|_{L^{\infty}(X)}$. By setting
$F_{j}((m_{1},\ldots,m_{k}))$ $=f_{j}(T_{1}^{m_{1}}\cdots T_{k}^{m_{k}}x),$
$G((m_{1},\ldots,m_{k}))$ $=\phi(T_{1}^{m_{1}}\cdots T_{k}^{m_{k}}x),$
the estimate (7.12) reduces to
$\left|\frac{1}{N^{k+1}}\sum_{\mathbf{m}\in[N]^{k},n\in[N]}\mu(n)G(\mathbf{m})F_{1}(\mathbf{m}+(n,0,\ldots,0))\cdots F_{k}(\mathbf{m}+(0,\ldots,0,n))\right|$
$\quad\ll_{A}\|F_{1}\|_{L^{\infty}(X)}\cdots\|F_{k}\|_{L^{\infty}(X)}\|G\|_{L^{\infty}(X)}(\log N)^{-A}.$
But this bound follows directly from Lemmas 4.7 and 5.1. ∎